OpenStudy (laddiusmaximus):

I need walked through this question: take the derivative of y=log2 3x

OpenStudy (amistre64):

base 2 each side

OpenStudy (amistre64):

2^y = 3x now solve as usual \[[B^x]'=B^x~ln(B)\]

OpenStudy (amistre64):

2^y ln(2) y' = 3 y' = 3/(2^y ln(2)) and since y = L.2(3x) y' = 3/(2^(L.2(3x)) ln(2)) y' = 3/(3x ln(2)) y' = 1/(x ln(2))

OpenStudy (laddiusmaximus):

wait multiply two to both sides?

OpenStudy (amistre64):

no, BASE 2 each side

OpenStudy (amistre64):

you have a log, you can modify it to an exponent if need be

OpenStudy (amistre64):

otherwise change of base the log and run the ln(3x) to fruition

OpenStudy (amistre64):

\[L_b(a)=\frac{ln(a)}{ln(b)}\]\[L_2(3x)=\frac{ln(3x)}{ln(2)}\]

OpenStudy (laddiusmaximus):

the answer is log2e/x

OpenStudy (laddiusmaximus):

how do we get there?

OpenStudy (amistre64):

thats not "the" answer. that is just the answer they chose to "simplify" to.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=d%2Fdx+log_2%283x%29 how they got to your answer, is anyones guess

OpenStudy (laddiusmaximus):

gotcha, thanks.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1%2F%28xln%282%29%29-log_2%28e%29%2Fx just so you know, the 2 are equal :)

OpenStudy (anonymous):

To get to the answer in the book: \[y' = \frac{1}{x \ln 2} = \frac{\ln e}{x \ln{2}}\] Since \(\log_2 e = {\ln e \over \ln 2}\) \[\frac{1}{x}\frac{\ln e}{\ln 2} = \frac{\log_2 e}{x}\]

OpenStudy (laddiusmaximus):

not really seeing it at the moment, but ill do a few more to try and understand it. that, or just move on to intergrals.